# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ 5 Nonsmooth Stochastic Implicit Differentiation
In this section we study the stochastic fixed point formulation in (2) and present an algorithm that, given a random vector $y\in\mathbb{R}^{d}$ and an approximate solution $w_{t}(\lambda)$, efficiently approximates an element of $D_{w}^{\operatorname{imp}}(\lambda)^{\top}y$ accessing only $\hat{T}$, $G$ and their conservative derivatives. Similarly to the deterministic case of AID, here we assume that $w_{t}(\lambda)$ is generated by a stochastic algorithm which converges in mean square to $w(\lambda)$. Several algorithms can ensure such convergence guarantees for the composite minimization problems in (7) (e.g, Rosasco et al. (2020) provide a proximal stochastic gradient algorithm with rate $O(1/t)$) and composite monotone inclusions (Rosasco et al., 2014).
@@
    $\hat{T}\colon\mathbb{R}^{d}\times O_{\Lambda}\times\Xi\to\mathbb{R}^{d}$, $\xi$ is a random variable with value in $\Xi$ and $\mathbb{E}[\hat{T}(u,\lambda,\xi)]=T(u,\lambda)$.
@@
    For every $x\in\Xi$, $\hat{T}(\cdot,\cdot,x)$ is path differentiable and
    $\partial_{1}\hat{T}\colon\mathbb{R}^{d}\times O_{\Lambda}\times\Xi\to\mathbb{R}^{d\times d},\ \ \partial_{1}\hat{T}\colon\mathbb{R}^{d}\times O_{\Lambda}\times\Xi\to\mathbb{R}^{d\times p}$
    are selections of $D_{\hat{T},1}$ and $D_{\hat{T},2}$ respectively and there exist $\sigma_{1},\sigma_{2},\sigma^{\prime}_{1},\sigma^{\prime}_{2}\geq 0$ such that for every $u\in\mathbb{R}^{d}$ and $\lambda\in\Lambda$
    $\mathbb{E}[\partial_{1}\hat{T}(u,\lambda,\xi)]=\partial_{1}T(u,\lambda)\in D_{T,1}(u,\lambda),\ \mathbb{E}[\partial_{2}\hat{T}(u,\lambda,\xi)]=\partial_{2}T(u,\lambda)\in D_{T,2}(u,\lambda),$
    $\operatorname{Var}[\hat{T}(u,\lambda,\xi)]\leq\sigma_{1}+\sigma_{2}\lVert u-T(u,\lambda)\rVert^{2},\quad\operatorname{Var}[\partial_{1}\hat{T}(u,\lambda,\xi)]\leq\sigma^{\prime}_{1},\quad\operatorname{Var}[\partial_{2}\hat{T}(u,\lambda,\xi)]\leq\sigma^{\prime}_{2}.$
In view of the chain rule for conservative derivatives 5.1 ensures that $\Phi$ is path differentiable and $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$. Thus, $w(\lambda)$ is well defined and it has conservative derivatives $D_{w}^{\operatorname{imp}}$ and $D_{w}^{\operatorname{fix}}$. 5.2 is a nonsmooth generalization of the corresponding one in (Grazzi et al., 2021, 2023). Finally, recalling (10), we note that
@@
Our method is inspired by (17) and (18) but it uses mini-batch estimators of $T$ and $\partial_{2}\Phi$. To that purpose we assume to have two independent sets of samples $\bm{\xi}^{(1)}=(\xi^{(1)}_{j})_{1\leq j\leq J}$ and $\bm{\xi}^{(2)}=(\xi^{(2)}_{i})_{1\leq i\leq k}$, being i.i.d. copies of the random variable $\xi$. Moreover, we define
$\bar{T}(u,\lambda){=}\frac{1}{J}\sum_{j=1}^{J}\hat{T}(u,\lambda,\xi^{(1)}_{j}),\quad\bar{\Phi}(u,\lambda)=G(\bar{T}(u,\lambda),\lambda).$
The function $\bar{\Phi}$ is nonsmooth and we can define its partial conservative derivatives as
$D_{\bar{\Phi},1}(u,\lambda)$ $=\!D_{G,1}(\bar{T}(u,\lambda),\lambda)D_{\bar{T},1}(u,\lambda)$
$D_{\bar{\Phi},2}(u,\lambda)$ $=\!D_{G,1}(\bar{T}(u,\lambda),\lambda)D_{\bar{T},2}(u,\lambda)\!+\!D_{G,2}(\bar{T}(u,\lambda)).$
In fact our approach replaces the linear system (18) with
@@
where the solution is in turn approximated by a stochastic sequence $(v_{k})_{k\in\mathbb{N}}$, which has access only to $\hat{T},G$, and $w_{t}(\lambda)$.
We first provide a general bound for the mean square error of an estimator of an element of the Jacobian vector product $D_{w}^{\operatorname{imp}}(\lambda)^{\top}y$, which is agnostic with respect to the algorithms solving the fixed point equation (1) and the linear system (19). The proof (in Appendix D) uses similar techniques as the one for the smooth case in (Grazzi et al., 2021, 2023).
@@
Under 5.1, 5.2, and 5.3, let $\kappa=(1-q)^{-1}$. We define the estimator
$(w^{\prime}(\lambda)^{\top}y)^{\wedge}:=\partial_{2}\bar{\Phi}(w_{t}(\lambda),\lambda)^{\top}v_{k}(w_{t}(\lambda),\bar{T}(w_{t}(\lambda),\lambda)).$
@@
$\mathbb{E}\big{[}\!\operatorname{\it{e}}\!\big{(}(w^{\prime}(\lambda)^{\top}y)^{\wedge},D_{w}^{\operatorname{imp}}(\lambda)^{\top}y\big{)}^{2}\big{]}=B^{2}\times O\left(\sigma_{\lambda}(k)+\kappa^{4}\left(\frac{1}{J}+\rho_{\lambda}(t)\right)\right).$
We preset the full procedure, named nonsmooth stochastic implicit differentiation (NSID), in Algorithm 1, where the sequence $v_{k}$ considered in Assumption 5.3(ii) is generated by a simple stochastic fixed-point iteration algorithm (described in (Grazzi et al., 2021)) with step sizes $(\eta_{i})_{1\leq i\leq k}$.
1:  Input: $k,J\in\mathbb{N}$, $w_{t}(\lambda),y\in\mathbb{R}^{d}$, $\hat{T}$, $G$, $\bm{\xi}^{(1)},\bm{\xi}^{(2)}$
@@
3:  $\hat{\Psi}\colon(v,x)\mapsto\partial_{1}\hat{T}(w_{t}(\lambda),\lambda,x)^{\top}\partial_{1}G(\bar{T}_{t}(\lambda),\lambda)^{\top}v+y$
@@
7:  Return $(w^{\prime}(\lambda)^{\top}y)^{\wedge}:=\partial_{2}\bar{\Phi}(w_{t}(\lambda),\lambda)^{\top}v_{k}$
@@
Note that all steps can be efficiently implemented via automatic differentiation by using only vector-valued function evaluations and conservative Jacobian-vector products without the expensive computation of the full matrix derivatives. Furthermore, using a fixed selection for the conservative derivative corresponds to the standard implementation.
If $G(\cdot,\lambda)$ is the identity and $T$ is smooth, we recover exactly the same procedure given in (Grazzi et al., 2021, 2023). We stress that handling a general $G$ provides an additional challenge since we do not have access anymore to an unbiased estimator of $\Phi$. However, we could overcome this issue by using different samples sequences for the two factors occuring in $\hat{\Psi}$. Incidentally, one of those can be the one used to compute a mini-batch estimator of $\partial\Phi_{2}$. Ultimately, this does not call for any additional samples compared to the smooth version.
@@
Under 5.1, 5.2, and 5.3(i)(iii), let $(w^{\prime}(\lambda)^{\top}y)^{\wedge}$ be generated by Algorithm 1 with $\eta_{i}=\Theta(i^{-1})$ and assume that $\rho_{\lambda}(t)=O(\kappa^{\alpha}t^{-1})$, with $\alpha>0$. Then
$\begin{split}\mathbb{E}\big{[}\!\operatorname{\it{e}}\!\big{(}(w^{\prime}(\lambda)^{\top}y)^{\wedge},D_{w}^{\operatorname{imp}}(\lambda)^{\top}y\big{)}^{2}\big{]}=O\left(\frac{\kappa^{5}}{k}+\frac{\kappa^{4}}{J}+\frac{\kappa^{4+\alpha}}{t}\right).\end{split}$
@@
Note that the sample complexity $O(\epsilon^{-1})$ matches the performance of SGD for minimizing strongly convex and Lipschitz smooth functions (Bottou et al., 2018), which are a special cases of Problem (2). Furthermore it is the same one that the SID algorithm by Grazzi et al. (2020, 2023) attains when $G(v,\lambda)=v$ and $\Phi$ is Lipschitz smooth. A limitation is the choice of step-sizes $(\eta_{i})$, problematic in practice.
@@ 6 Application to Bilevel Optimization
First, note that if $E$ and $\Phi$ are path differentiable with conservative derivatives $D_{E}$ and $D_{\Phi}$ respectively and if $\lVert D_{\Phi}(u,\lambda)\rVert_{\sup}\leq q<1$ for every $u\in\mathbb{R}^{d}$, $\lambda\in\Lambda$, then
$D_{f}^{\operatorname{imp}}(\lambda)$ ${:=}D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,D_{E,1}(w(\lambda),\lambda){+}D_{E,2}(w(\lambda),\lambda)$
$D_{f}^{\operatorname{fix}}(\lambda)$ ${:=}D_{w}^{\operatorname{fix}}(\lambda)^{\top}D_{E,1}(w(\lambda),\lambda){+}D_{E,2}(w(\lambda),\lambda)$
are conservative derivatives for $f$. We further assume that $E$ satisfies Assumption 3.1(i) with corresponding derivative $D_{E}$. Let also set $f_{t}(\lambda):=E(w_{t}(\lambda),\lambda)$, where $w_{t}(\lambda)$ is an approximate solution for the fixed point problem. In the following we let $z_{t}=(w_{t}(\lambda),\lambda)$ for brevity.
@@
Bilevel AID-FP (BAID-FP) and Bilevel ITD (BITD) approximate $D_{f}$ by setting $t,k\in\mathbb{N}$, and returning an element of the sets
@@
This guarantees the same convergence rates, this time to $D_{f}^{\operatorname{fix}}(z)$, as ITD and AID (Theorem E.3).
@@
In the stochastic case, the bilevel problem takes the form
$\begin{gathered}\min_{\lambda\in\Lambda}f(\lambda):=\mathbb{E}[\hat{E}(w(\lambda),\lambda,\zeta)],\\ w(\lambda)=G\big{(}\mathbb{E}[\hat{T}(w(\lambda),\lambda,\xi)],\lambda\big{)}.\end{gathered}$ (21)
where $\zeta$ is a random variable. We consider Algorithm 2, which additionally computes a minibatch gradient estimator of $E^{\prime}\in D_{E}$, using the sequence $\bm{\zeta}^{(1)}=(\zeta^{(1)})_{1\leq i\leq J_{1}}$ of i.i.d. copies of $\zeta$.
1:  Input: $k,J_{1},J_{2}$, $w_{t}(\lambda)$, $\hat{E}$, $\hat{T}$, $G$, $\bm{\xi}^{(1)},\bm{\xi}^{(2)},\bm{\zeta}^{(1)}$
2:  $\bar{E}^{\prime}(z_{t})\leftarrow J_{1}^{-1}\sum_{i=1}^{J_{1}}\hat{E}^{\prime}(z_{t},\zeta^{(1)}_{i})$
3:  $r(z_{t}){\leftarrow}\mathrm{NSID}(k,J_{2},w_{t}(\lambda),\partial_{1}\bar{E}(z_{t}),\hat{T},G,\bm{\xi}^{(1)},\bm{\xi}^{(2)})$
4:  Return ^$\hat{\nabla}f(\lambda):=r(z_{t})+\partial_{2}\bar{E}(z_{t})$
@@
Despite these encouraging results and the fact that in the smooth case several works provide convergence rates to a stationary point of the gradient of $f$ (Ji et al., 2021; Arbel & Mairal, 2021; Grazzi et al., 2023), proving such type of results or even asymptotic convergence (without rates) in our nonsmooth case is more challenging and we leave it for future work. One crucial issue is that in the analysis, the constant defined in Lemma 2.5, which we use in place of that of Lipschitz smoothness, cannot be properly controlled on the whole $\Lambda$ as required in the smooth case: it becomes arbitrarily large when $(w(\lambda),\lambda)$ approaches non-differentiable regions of $\Phi$.

### v2

@@ 5 Nonsmooth Stochastic Implicit Differentiation
In this section we study the stochastic fixed point formulation in (2) and present an algorithm that, given a random vector $y\in\mathbb{R}^{d}$ and an approximate solution $w_{t}(\lambda)$, efficiently approximates an element of $D_{w}^{\operatorname{imp}}(\lambda)^{\top}y$ accessing only $\hat{T}_{\xi}$, $G$ and fixed selections of their conservative derivatives. Similarly to deterministic AID, here we assume that $w_{t}(\lambda)$ is generated by a stochastic algorithm which converges in mean square to $w(\lambda)$. Several algorithms can ensure such convergence guarantees for the composite minimization problems in (8) (e.g, Rosasco et al. (2020) provide a proximal stochastic gradient algorithm with rate $O(1/t)$) and composite monotone inclusions (Rosasco et al., 2014).
We recall that for a path differentiable function $F\colon U\subset\mathbb{R}^{p_{1}+p_{2}}\to\mathbb{R}^{d}$, we denote by $F^{\prime}$ an arbitrary selection of $D_{F}$ and by $\partial_{1}F(x)\in\mathbb{R}^{d\times p_{1}}$ and $\partial_{2}F(x)\in\mathbb{R}^{d\times p_{2}}$ the first and second block component of $F^{\prime}(x)$ respectively, so that we can write $F^{\prime}(x)=[\partial_{1}F(x),\partial_{2}F(x)]$.
@@
    $\hat{T}_{x}\colon\mathbb{R}^{d}\times O_{\Lambda}\to\mathbb{R}^{d}$ and $\mathbb{E}[\hat{T}_{\xi}(u,\lambda)]=T(u,\lambda)$.
@@
    $\hat{T}_{x}$ is path differentiable and $\hat{T}_{x}^{\prime}$ is a selection of its conservative derivative $D_{\hat{T}_{x}}$ and there exist $\sigma_{1},\sigma_{2},\sigma^{\prime}_{1},\sigma^{\prime}_{2}\geq 0$ such that for every $u\in\mathbb{R}^{d}$, $\lambda\in\Lambda$
    $\mathbb{E}[\hat{T}^{\prime}_{\xi}(u,\lambda)]=T^{\prime}(u,\lambda)\in D_{T}(u,\lambda),$
    $\operatorname{Var}[\hat{T}_{\xi}(u,\lambda)]\leq\sigma_{1}+\sigma_{2}\lVert u-T(u,\lambda)\rVert^{2},\quad\operatorname{Var}[\partial_{1}\hat{T}_{\xi}(u,\lambda)]\leq\sigma^{\prime}_{1},\quad\operatorname{Var}[\partial_{2}\hat{T}_{\xi}(u,\lambda)]\leq\sigma^{\prime}_{2}.$
    where $\hat{T}^{\prime}_{x}(u,\lambda)=[\partial_{1}\hat{T}_{x}(u,\lambda),\partial_{2}\hat{T}_{x}(u,\lambda)]$.
Remark 5.3.
The above assumptions can be satisfied in the following situations: (1) $G$ is nonsmooth, e.g., some proximity operator or the projection on some simple constraints, while $T$ and $\hat{T}_{x}$ are smooth (e.g., one step of gradient descent of a twice differentiable loss); (2) in view of Remark 3.3, when $T=\frac{1}{n}\sum_{i=1}^{n}\hat{T}_{i}$, $T^{\prime}=\frac{1}{n}\sum_{i=1}^{n}\hat{T}^{\prime}_{i}$ with $\hat{T}^{\prime}\in D_{\hat{T}_{i}}^{s}$ and $\xi$ is uniformly distributed on $[n]$.
5.1 ensures that $D_{\Phi}$ obtained via the chain rule for conservative derivatives in (Bolte & Pauwels, 2021) (see Appendix D) is a conservative derivative of $\Phi$ and that $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$. Thus, $w(\lambda)$ is well defined and it has conservative derivatives $D_{w}^{\operatorname{imp}}$ and $D_{w}^{\operatorname{fix}}$. 5.2 is a nonsmooth generalization of the corresponding one in (Grazzi et al., 2021, 2023). Finally, recalling (11), if $\partial_{2}\Phi(u,\lambda)=\partial_{1}G(T(u,\lambda),\lambda)\partial_{2}T(u,\lambda)+\partial_{2}G(T(u,\lambda),\lambda)$ then
@@
Our method is inspired by (18) and (19) but it uses mini-batch estimators of $T$ and $\partial_{2}\Phi$. To that purpose we assume to have two independent sets of samples $\bm{\xi}^{(1)}=(\xi^{(1)}_{j})_{1\leq j\leq J}$ and $\bm{\xi}^{(2)}=(\xi^{(2)}_{i})_{1\leq i\leq k}$, being i.i.d. copies of the random variable $\xi$. Moreover, we define the path differentiable functions
$\bar{T}(u,\lambda){=}\frac{1}{J}\sum_{j=1}^{J}\hat{T}_{\xi^{(1)}_{j}}(u,\lambda),\quad\bar{\Phi}(u,\lambda)=G(\bar{T}(u,\lambda),\lambda).$
In fact our approach first replaces the linear system (19) with
@@
where the solution is in turn approximated by a stochastic sequence $(v_{k})_{k\in\mathbb{N}}$, which has access only to $\hat{T}_{x},G$, and $w_{t}(\lambda)$. Second, it outputs $\partial_{2}\bar{\Phi}(w_{t}(\lambda),\lambda)^{\top}v_{k}$, where for any $u\in\mathbb{R}^{d}$, $\lambda\in O_{\Lambda}$,
$\partial_{2}\bar{\Phi}(u,\lambda)$ ${=}\partial_{1}G(\bar{T}(u,\lambda),\lambda)\partial_{2}\bar{T}(u,\lambda)+\partial_{2}G(\bar{T}(u,\lambda),\lambda),$ (21)
with $\bar{T}^{\prime}(u,\lambda){:=}[\partial_{1}\bar{T}(u,\lambda),\partial_{1}\bar{T}(u,\lambda)]{=}\frac{1}{J}\sum_{j=1}^{J}\hat{T}^{\prime}_{\xi_{j}^{(1)}}(u,\lambda)$, which thanks to the chain rule is an element of a partial conservative derivative of $\bar{\Phi}$ (see also Appendix D).
We now provide a general bound for the mean square error of an estimator of an element of the Jacobian vector product $D_{w}^{\operatorname{imp}}(\lambda)^{\top}y$, which is agnostic with respect to the algorithms solving the fixed point equation (1) and the linear system (20). The proof (in Appendix D) uses similar techniques as the one for the smooth case in (Grazzi et al., 2021, 2023).
@@
Under 5.1, 5.2, and 5.4, let $\kappa=(1-q)^{-1}$. We define the estimator
$(w^{\prime}(\lambda)^{\top}y)^{\widehat{}}:=\partial_{2}\bar{\Phi}(w_{t}(\lambda),\lambda)^{\top}v_{k}\big{(}w_{t}(\lambda),\bar{T}(w_{t}(\lambda),\lambda)\big{)}.$
@@
$\mathbb{E}\big{[}\!\operatorname{\it{e}}\!\big{(}(w^{\prime}(\lambda)^{\top}y)^{\widehat{}},$ $D_{w}^{\operatorname{imp}}(\lambda)^{\top}y\big{)}^{2}\big{]}=b^{2}\times O\left(\sigma_{\lambda}(k)+\kappa^{4}\left(J^{-1}+\rho_{\lambda}(t)\right)\right).$
We preset the full procedure, named nonsmooth stochastic implicit differentiation (NSID), in Algorithm 1, where the sequence $v_{k}$ considered in Assumption 5.4(ii) is generated by a simple stochastic fixed-point iteration algorithm (described in (Grazzi et al., 2021) and recalled in Appendix D) with step sizes $(\eta_{i})_{1\leq i\leq k}$.
1:  Input: $k,J\in\mathbb{N}$, $w_{t}(\lambda),y\in\mathbb{R}^{d}$, $\bm{\xi}^{(1)},\bm{\xi}^{(2)}$
@@
3:  $\hat{\Psi}\colon(v,x)\mapsto\partial_{1}\hat{T}_{x}(w_{t}(\lambda),\lambda)^{\top}\partial_{1}G(\bar{T}_{t}(\lambda),\lambda)^{\top}v+y$
@@
7:  Return $(w^{\prime}(\lambda)^{\top}y)^{\widehat{}}:=\partial_{2}\bar{\Phi}(w_{t}(\lambda),\lambda)^{\top}v_{k}$
@@
Note that all steps can be efficiently implemented via automatic differentiation by using only vector-valued function evaluations and conservative Jacobian-vector products without the expensive computation of the full matrix derivatives. Also, using a fixed selection for the conservative derivative of $\hat{T}_{x}$ and $G$ corresponds to the standard implementation.
If $G(\cdot,\lambda)$ is the identity and $T$ is smooth, NSID reduces to the same procedure given in (Grazzi et al., 2023), which also provide the bound $O\left(\sigma_{\lambda}(k)+\kappa^{2}J^{-1}+\kappa^{4}\rho_{\lambda}(t)\right)$ in Theorem 7. Compared to the bound given in Theorem 5.5, we note that the only difference is in the constant in front of the term $J^{-1}$, which we believe may be related to the term $G$. Indeed handling a general $G$ provides an additional challenge since we do not have access anymore to an unbiased estimator of $\Phi$. However, we could overcome this issue by using different samples sequences for the two factors occuring in $\hat{\Psi}$. Incidentally, one of those can be the one used to compute a mini-batch estimator of $\partial_{2}\Phi$. Ultimately, this does not call for any additional samples compared to the smooth version, but it could worsen some constants in the bound.
@@
Under 5.1, 5.2, and 5.4(i)(iii), let $(w^{\prime}(\lambda)^{\top}y)^{\widehat{}}$ be generated by Algorithm 1 with $\eta_{i}=\Theta(i^{-1})$ and assume that $\rho_{\lambda}(t)=O(\kappa^{\alpha}t^{-1})$, with $\alpha>0$. Then
$\begin{split}\mathbb{E}\big{[}\!\operatorname{\it{e}}\!\big{(}(w^{\prime}(\lambda)^{\top}y)^{\widehat{}},D_{w}^{\operatorname{imp}}(\lambda)^{\top}y\big{)}^{2}\big{]}=O\left(\frac{\kappa^{5}}{k}{+}\frac{\kappa^{4}}{J}{+}\frac{\kappa^{4+\alpha}}{t}\right).\end{split}$
@@
Note that the sample complexity $O(\epsilon^{-1})$ matches the performance of SGD for minimizing strongly convex and Lipschitz smooth functions (Bottou et al., 2018), which are a special cases of Problem (2). Furthermore it is the same one that the SID algorithm by Grazzi et al. (2021, 2023) attains when $G(v,\lambda)=v$ and $\Phi$ is Lipschitz smooth. A limitation is the choice of step-sizes $(\eta_{i})$, problematic in practice.
@@ 6 Application to Bilevel Optimization
In addition to the requirement that $\Phi$ satisfies Assumption 3.1, we also make the hypothesis that $E$ satisfies the first item of same assumption with corresponding conservative derivative $D_{E}=D_{E}^{s}$. Therefore, applying the usual chain rule, we have that for $*\in\{\operatorname{imp},\operatorname{fix}\}$
$D_{f}^{*}(\lambda)$ ${:=}D_{E}(w(\lambda),\lambda)\,\begin{bmatrix}D_{w}^{*}(\lambda)\\ I_{m}\end{bmatrix}$
is a conservative derivatives for $f$. We also let $f_{t}(\lambda):=E(w_{t}(\lambda),\lambda)$, where $w_{t}(\lambda)$ is an approximate solution for the fixed point problem.
@@
The approximate derivatives
@@
converge to $D_{f}^{\operatorname{fix}}(\lambda)$ with the same rate as ITD and AID (Theorem E.3).
@@
We study the bilevel problem
$\begin{gathered}\min_{\lambda\in\Lambda}f(\lambda):=\mathbb{E}[\hat{E}_{\zeta}(w(\lambda),\lambda)],\\ w(\lambda)=G\big{(}\mathbb{E}[\hat{T}_{\xi}(w(\lambda),\lambda)],\lambda\big{)}.\end{gathered}$ (23)
where $\zeta$ is a random variable. We consider Algorithm 2, which additionally computes $\bar{E}^{\prime}(w_{t}(\lambda),\lambda):=J_{1}^{-1}\sum_{j=1}^{J_{1}}\hat{E}_{\zeta^{(1)}_{j}}^{\prime}(w_{t}(\lambda),\lambda)$, a minibatch gradient estimator of $E^{\prime}\in D_{E}$, using the sequence $\bm{\zeta}^{(1)}=(\zeta^{(1)})_{1\leq j\leq J_{1}}$ of i.i.d. copies of $\zeta$.
1:  Input: $k,J_{1},J_{2}\in\mathbb{N}$ , $w_{t}(\lambda)\in\mathbb{R}^{d}$, $\bm{\xi}^{(1)},\bm{\xi}^{(2)},\bm{\zeta}^{(1)}$
2:  Compute $\bar{E}^{\prime}(w_{t}(\lambda),\lambda)$        (using $\bm{\zeta}^{(1)}$)
3:  $y\leftarrow\partial_{1}\bar{E}(w_{t}(\lambda),\lambda)^{\top}$
4:  $r(w_{t}(\lambda),\lambda)\leftarrow\mathrm{NSID}(k,J_{2},w_{t}(\lambda),y,\bm{\xi}^{(1)},\bm{\xi}^{(2)})$
5:  Return ^$\hat{\nabla}f(\lambda)^{\top}:=r(w_{t}(\lambda),\lambda)^{\top}+\partial_{2}\bar{E}(w_{t}(\lambda),\lambda)$
@@
Despite these encouraging results and the fact that in the smooth case several works provide convergence rates to a stationary point of the gradient of $f$ (Ji et al., 2021; Arbel & Mairal, 2021; Grazzi et al., 2023), proving such type of results or even asymptotic convergence (without rates) in our nonsmooth case is more challenging and we leave it for future work. One crucial issue is that in the analysis, the constant defined in Lemma 2.5, which we use in place of that of Lipschitz smoothness, cannot be properly controlled on the whole $\Lambda$ as required in the smooth case: it becomes arbitrarily large when $(w(\lambda),\lambda)$ approaches nondifferentiable regions of $\Phi$.